OpenStudy (anonymous):

What are the next 3 terms of the following sequence? 3, -1, -5, -9

OpenStudy (anonymous):

a=3 ,d=-4 you have two ways : 1st way to add -4 to the last term and so on ... 2nd way to use the rule Tn=a+(n-1)d t5=3+4*-4=-13 t6=3+5*-4=-17 t7=3+6*-4=-21 Got it :D

OpenStudy (anonymous):

THANKSS FOR ALL YOU R HELP!

OpenStudy (anonymous):

Your very welcomeee xD